### v2

# Generative Refocusing: Flexible Defocus Control from a Single Image

Abstract
Depth-of-field control is essential in photography, but getting the perfect focus often takes several tries or special equipment. Single-image refocusing is still difficult. It involves recovering sharp content and creating realistic bokeh. Current methods have significant drawbacks. They need all-in-focus inputs, depend on synthetic data from simulators, and have limited control over aperture. We introduce Generative Refocusing, a two-step process that uses DeblurNet to recover all-in-focus images from various inputs and BokehNet for creating controllable bokeh. Our main innovation is semi-supervised training. This method combines synthetic paired data with unpaired real bokeh images, using EXIF metadata to capture real optical characteristics beyond what simulators can provide. Our experiments show we achieve top performance in defocus deblurring, bokeh synthesis, and refocusing benchmarks. Additionally, our Generative Refocusing allows text-guided adjustments and custom aperture shapes. Project page: https://generative-refocusing.github.io/
1 Introduction
Controlling depth of field and focus is essential for artistic photography. It helps guide the viewer’s eye through selective focus and bokeh effects. Getting perfect focus often requires multiple shots or special equipment, which can be tough for everyday photographers. Single-image refocusing removes these obstacles. It allows for focus and bokeh adjustments after the photo is taken. The main challenge is recovering sharp details from blurred areas while creating natural-looking bokeh, while still giving users fine control.
Existing research addresses different aspects of this problem. Defocus deblurring methods [1, 8, 36, 56, 67, 73, 84] aim to recover sharp images. Diffusion approaches [41, 42] show promise for addressing spatially variant blur. Bokeh synthesis methods [9, 30, 48, 60] focus on rendering realistic depth-of-field. Recent efforts include physics-based [38, 37, 50], neural [48, 49], and diffusion-based [19, 82, 71] approaches. However, most methods handle either deblurring or bokeh synthesis individually. End-to-end refocusing often requires specialized capture setups [40, 46, 2].
Despite these advances, three main limitations remain (Tab. 1). First, most methods rely on all-in-focus inputs and accurate depth maps. This restricts their use to images that already have defocus blur. Second, while synthetic data [48, 89] can be used for training, it is limited by the realism of the simulators. This approach often fails to capture detailed visual appearances. On the other hand, obtaining real paired datasets [60] is very expensive. Third, current methods usually only support aperture size, not its shape. This restriction limits the possibilities for creative bokeh effects.
Table 1: Comparison across methods.  = supported;  = not $\LARGE\triangle = supported$ with caveat. In the synthetic/real training data columns, “paired” indicates the use of paired supervision, while “unpaired” indicates real unpaired data (no real paired supervision). Bokehlicious typically assumes an all-in-focus input and requires fine-tuning for different input types. For BokehMe, only the classical renderer variant supports aperture-shape control. UW = ultra-wide, W = wide.
Method Input Training data Supported user control
Setup AIF required? Synthetic Real Aperture size Aperture shape Focus point Defocus map
[…]
DC2 [2] Dual-view (UW+W) paired
Bokehlicious [60] Single-view paired
BokehDiff [89] Single-view paired
DiffCamera [71] Single-view paired
Ours Single-view paired unpaired
To tackle these issues, we present Generative Refocusing, a flexible single-image refocusing system built on a two-stage design. Our DeblurNet module produces a sharp, in-focus image from a variety of inputs. It uses a diffusion model guided by initial deblurring predictions. Our BokehNet creates fully customizable bokeh, considering user-defined focus planes, bokeh intensity, and aperture shapes (LABEL:fig:teaser). A significant advancement in our method is our training scheme, which makes use of both 1) paired synthetic data to maintain geometric consistency and 2) unpaired real-world bokeh images with camera metadata that help us learn real lens characteristics not captured by simulators. Our approach achieves top performance in defocus deblurring, bokeh generation, and refocusing, while maintaining the natural consistency of scenes.
Our main contributions are:
  • A flexible refocusing pipeline accepts images in any focus state. It gives users control over the focus plane, bokeh intensity, and aperture shape through two-stage decomposition.
  • This semi-supervised framework connects synthetic and unpaired real bokeh images. It enables genuine optical learning that goes beyond what a simulator can offer. This is achieved through parameter estimation from EXIF metadata.
  • The system performs well across defocus deblurring, bokeh synthesis, and refocusing. This has been validated on existing benchmarks (DPDD [1], RealDOF [36]) and our new light-field datasets (LF-Bokeh, LF-Refocus). It also has applications in text-guided deblurring and creating aperture shapes.
2 Related Work
Figure 2: Pipeline Overview. Our method decomposes single-image refocusing into two stages: (a) Defocus Deblurring and (b) Shallow Depth-of-Field Synthesis. Given a blurry input image $I_{\text{in}}$, we optionally apply a pre-deblurring method [56] to obtain a conservative estimate $I_{\text{syn}}$, then feed both $I_{\text{in}}$ and $I_{\text{syn}}$ into DeblurNet to recover a high-quality all-in-focus image $I_{\text{aif}}$. The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ convert images to latent representations for processing by the DiT backbone. In the second stage, BokehNet takes the all-in-focus image $I_{\text{aif}}$, the defocus map $D_{\text{def}}$, and optionally a specific aperture shape as inputs to synthesize the refocused output. The defocus map $D_{\text{def}}$ is computed based on the estimated depth map $D$ [7], along with both the user-specified focus plane $S_{1}$ and bokeh level $K$.
Diffusion Models for Image Restoration.
Diffusion models [24, 65, 81, 68] advance image restoration via generative priors, extending to zero-shot restoration and artifact removal. Methods evolved from pixel-space [57] to efficient latent-space approaches [54, 76], enabling faster inference. ResShift [83] cuts steps from 1000 to 4–15 via residual shifting. Deblurring progressed from merging diffusion priors with regression tasks, such as matting [74] or hierarchical restoration [13], to learning spatially-varying kernels in latent space [33]. Cascaded pipelines [25] refine at multiple resolutions, while principled frameworks [45, 44] treat degradations as stochastic processes. Training-free methods [91, 32] enable plug-and-play restoration. Unlike these general methods, we target spatially-varying defocus blur with a two-stage pipeline explicitly separating deblurring from bokeh synthesis for controllable refocusing.
Defocus Deblurring.
Defocus deblurring recovers sharp images from spatially varying blur caused by limited depth of field. Early deconvolution [4, 92] introduced artifacts at depth transitions. Dual-pixel sensors [1], quad-pixel data [8], and disparity-aware techniques [36, 79, 70] improved blur–depth modeling and the recovery of all-in-focus images from focal stacks. Architectures shifted from CNNs [10] to transformers [67, 73], with gains from implicit representations [51, 17] and multi-scale attention [87]. Recent work includes vision-language fusion [77] for semantic estimation and a shift from supervised [52] to diffusion-based learning [42, 41], including unpaired data. While these excel as standalone solutions, we position deblurring as the first stage of refocusing, using FLUX’s [61] generative prior to enable controllable bokeh synthesis from any input.
Bokeh Rendering.
Bokeh rendering progressed from physically based scattering [50, 38, 37] and differentiable rendering [63] to neural methods combining ray tracing with learned enhancements [48, 49]. Learning-based approaches evolved from fixed apertures [30] to variable f-stops [60, 9], extending to video [80] and 3D scenes [62, 72]. Diffusion models enabled bokeh control [19, 82], mainly for text-to-image synthesis. These methods require all-in-focus inputs and accurate depth. We remove this constraint via flexible input handling, which accepts defocused or all-in-focus images, and semi-supervised learning on unpaired real-world bokeh to capture authentic lens characteristics beyond simulators.
Single-Image Refocusing.
Post-capture refocusing manipulates focus after capture. Light-field cameras [46, 40] enable optical refocusing but need specialized capture. Computational approaches evolved from deconvolution [4, 86] producing artifacts, to GAN-based paradigms [58], to diffusion methods [71]. 3D representations [62, 72, 35, 43] achieve real-time controllable DoF but need multi-view capture or calibration. We combine three capabilities: (1) semi-supervised learning mixing synthetic and unpaired real data for authentic optics, (2) single-image flexibility accepting any input without preprocessing, and (3) comprehensive control over focus, intensity, and aperture shape.
Camera-Conditioned Diffusion.
Conditioning diffusion models on camera parameters progressed from extrinsic control for view synthesis [26, 14] and video generation [22, 75], to intrinsic control of focal length, aperture, and exposure [18, 12]. Direct conditioning often yields inconsistent content. Recent work uses temporal modeling [82], explicit calibration [6, 16], or transformer-based pose handling [3, 59]. These focus on generation; we operate in editing, manipulating existing images with scene consistency. By decoupling into deblurring and bokeh stages, we achieve efficient single-image manipulation without multi-frame overhead.
Semi-Supervised Learning.
When paired data is scarce, semi-supervised learning provides alternatives. The field evolved from learning without clean targets [39, 34, 27] to pseudo-labeling [11, 23, 69]. Foundation models enabled the generation of high-quality pseudo-labels [21, 85]. Unpaired frameworks [31, 90, 64] learn enhancement without paired supervision. Two-stage approaches [29, 78] bridge fidelity and perceptual quality via different supervision types. We adapt these to bokeh rendering: perfect paired refocusing data is nearly impossible to obtain. We train on synthetic pairs for structure, then unpaired real bokeh for realism. Deblur restores to all-in-focus, enabling bokeh to learn authentic lens characteristics exceeding simulators.
3 Method
Single-image refocusing couples two operations: recovering sharp content in out-of-focus regions and applying controllable bokeh to originally sharp areas. Blur magnitude is dictated by a defocus map parameterized by a user-specified focus point and the scene depth. When the input is blurry, monocular depth estimation is brittle, undermining precise control over the defocus map and often yielding misfocus or artifacts. We propose Generative Refocusing, which decomposes the task into defocus deblurring and bokeh synthesis, and comprises two models, DeblurNet and BokehNet, that jointly enable precise and controllable refocusing (see Fig. 2 for an overview of the pipeline); Such a flexible pipeline allows us to exert greater control over input conditions (see Tab. 1).
Figure 3: Training data generation. Each training sample consists of the following five components: (i) a bokeh image, (ii) an all-in-focus (AIF) image, (iii) a depth map $D$, (iv) a bokeh level $K$, and (v) a focus plane $S_{1}$. We construct these samples via three routes: (a) Synthetic paired data. Given real AIF images and depth maps $D$, we compute a defocus map $D_{\text{def}}$ parameterized by a specified bokeh level $K$ and focus plane $S_{1}$, and feed it into a bokeh renderer [48] to synthesize corresponding bokeh images. (b) Real unpaired data. Given real bokeh images, DeblurNet recovers an AIF image. We then estimate depth and extract a foreground mask [88] to define the focus plane $S_{1}$. The bokeh level $K$ is computed from the EXIF metadata following the formulation in [19]. (c) Real paired data without EXIF. For real pairs lacking EXIF, we obtain $S_{1}$ as in (b), and follow Eq. (2) to estimate the bokeh level $K$.
3.1 Stage 1: Defocus Deblurring
DeblurNet. We condition the deblurring network on a pair $(I_{\text{in}},I_{\text{pd}})$, where $I_{\text{pd}}$ is a conservative pre-deblur generated by an off-the-shelf defocus deblurring model. In our implementation, we use a non-generative defocus model  [56], whose outputs are typically content-faithful yet over-smoothed. Our model, guided by a learned diffusion prior, then injects high-fidelity details without sacrificing geometric correctness.
Artifact-robust dual conditioning.
To prevent artifacts in $I_{\text{pd}}$ from biasing the reconstruction and to ensure that fine details are recovered from $I_{\text{in}}$, we introduce two complementary mechanisms (see Fig. 2 (a)):
  • Positional disentanglement. We encode the two inputs with disjoint 2D positional grids: $I_{\mathrm{in}}$ is anchored at $(0,0)$, while $I_{\mathrm{pd}}$ is shifted to start at $(0,N)$, where $N$ denotes the side length of the token grid after downsampling $I_{\mathrm{in}}$. This shift lets the network distinguish the two sources, using $I_{\mathrm{pd}}$ as a low-frequency anchor while extracting high-frequency evidence from $I_{\mathrm{in}}$.
  • Pre-deblur dropout. During training, we stochastically replace $I_{\mathrm{pd}}$ with a zero tensor (at a fixed probability), forcing the network to remain effective when the auxiliary input is unreliable or absent. This regularizes the fusion, shields against method-specific artifacts, and enables the model to operate in a single-input deblurring mode at test time.
3.2 Stage 2: Shallow Depth-of-Field Synthesis
Synthetic-pair learning. BokehNet expects an AIF image, a target bokeh image, and the corresponding defocus map (see Fig. 2 (b)). In practice, perfectly paired data and accurate defocus maps are hard to acquire. We therefore pretrain with synthetic pairs: starting from real all-in-focus images and their estimated depth, we randomly sample a focus plane $S_{1}$ and a target bokeh level $K$, and use a simulator [48] to render bokeh consistent with the induced defocus map (see Fig. 3 (a)). This stage quickly trains the network to modulate the circle of confusion based on the defocus map, but it remains constrained by renderer bias and may introduce non-realistic artifacts.
Real unpaired bokeh learning. To capture real optics, we further leverage unpaired real bokeh images with EXIF metadata (see Fig. 3 (b)) and estimate an approximate bokeh level K\!K following [19]:
$K\;\approx\;\frac{f^{2}\,S_{1}}{\,2\,F\,(S_{1}-f)\,}\times\textit{pixel\_ratio},$ (1)
where $f$ is the focal length and $F$ denotes the aperture $f$-number; and pixel_ratio compensates for sensor size and image resolution differences. The AIF input is provided by DeblurNet. Although $S_{1}$ can sometimes be read from metadata, both estimated and true depths may deviate in real scenes. To obtain more reliable supervision, we estimate in-focus masks using BiRefNet [88], and have human experts filter erroneous cases and complete partial masks. These masks refine the focus plane and yield a consistent defocus map, enabling BokehNet to learn from real bokeh images.
Real paired bokeh images without metadata.
Previous bokeh-synthesis datasets [60] [30] provide real pairs but omit EXIF metadata or provide insufficient fields to estimate a reliable $K$ (see Fig. 3 (c)). We therefore adopt a simulator-in-the-loop calibration of the bokeh level: given an AIF image $I_{\mathrm{aif}}$, an estimated depth $D$, and a focus annotation obtained as in the unpaired setting (BiRefNet-based in-focus masks with expert filtering), we sweep $K$ and choose the value whose rendered result best matches the real bokeh target $I_{\mathrm{real}}$:
$K^{\star}=\operatorname*{argmax}_{\begin{subarray}{c}K\in\\ (K_{\min},\,K_{\max})\end{subarray}}\operatorname{SSIM}\!\big(B(I_{\mathrm{aif}},D;\,S_{1},K),\,I_{\mathrm{real}}\big),$ (2)
where B denotes our physically guided renderer and $S_{1}$ is the focus-plane proxy. The selected $K^{\star}$ is then used as the pseudo bokeh level label for training . Focus-plane annotation follows the same protocol as in the unpaired case.
Discussion.
Although the resulting defocus maps and the AIF proxy are not perfectly accurate, this supervision regime mirrors the practical refocusing scenario. Aggregating numbers of real bokeh images substantially improves realism and robustness over synthetic-only training.
3.3 Bokeh-Shape Aware Synthesis
State-of-the-art learning-based bokeh synthesis methods deliver compelling results but do not expose aperture-shape control. Physics-based renderers, in principle, support arbitrary shapes, yet public implementations typically omit this functionality. We therefore explore explicit shape-aware control using only a bokeh shape image within a trainable framework.
Table 2: Defocus deblurring benchmark on RealDOF [36] and DPDD [1] datasets. We evaluate the task of recovering all-in-focus images from defocused-blur inputs. Our method, GenRefocus, uses the DeblurNet module for this task. We report both reference-based metrics (LPIPS, FID) and no-reference perceptual quality metrics (CLIP-IQA, MANIQA, MUSIQ). Best, second best, and third best results are highlighted. GenRefocus outperforms all competing approaches across both benchmarks and all metrics, demonstrating the effectiveness of our diffusion-based deblurring with pre-deblur conditioning.
RealDOF [36] DPDD [1]
Method LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
Input 0.5241 92.93 0.3562 0.2213 28.7087 0.3485 88.85 0.4337 0.3325 45.5376
GT 0.4233 0.2675 42.5777 0.4563 0.3294 47.8102
AIFNet [55] 0.4132 72.38 0.3814 0.2536 35.7121 0.3090 79.84 0.4359 0.3371 46.9558
IFANet [36] 0.2994 33.21 0.4088 0.2609 39.5949 0.2301 42.60 0.4472 0.3276 47.2784
DRBNet [56] 0.2550 29.89 0.3889 0.2609 38.9014 0.1819 39.37 0.4142 0.3254 46.7229
Restormer [84] 0.2863 29.80 0.4062 0.2642 40.8355 0.1762 33.45 0.4332 0.3307 47.9872
INIKNet [51] 0.2851 32.18 0.3984 0.2580 38.9959 0.1860 36.63 0.4290 0.3265 47.1961
GenRefocus (Ours) 0.2356 24.73 0.4575 0.2886 43.7042 0.1598 33.08 0.4619 0.3411 49.3019
Figure 4: Qualitative comparison on defocus deblurring. Visual results on (a) RealDOF [36] and (b) DPDD [1] datasets. Blue boxes on the left indicate cropped regions shown in detail. Our DeblurNet faithfully recovers fine text details (“NEW YORK”) in (a) where competing methods produce blurry or distorted results. In the challenging example (b) with severe defocus blur, other methods either fail to recover structure (AIFNet [55], DRBNet [56], INIKNet [51]) or introduce artifacts in the background (IFANet [36], Restormer [84]), while our method reconstructs geometrically consistent, visually compelling content by leveraging diffusion priors guided by the pre-deblurred input.
Simulator and Data.
Real photographs exhibiting diverse bokeh shapes are rare, and paired AIF–bokeh examples are even scarcer. We thus rely on simulation. A key observation is that when an all-in-focus (AIF) image lacks point-light stimuli, the simulated bokeh carries weak shape cues, making the model reluctant to learn shape-conditioned responses. To address this, we synthesize a PointLight-1K AIF set designed to reveal aperture shape. We also extend the classical BokehMe renderer to scatter through a given shape: given an AIF image $I_{\mathrm{aif}}$, depth $D$, focus-plane proxy $S_{1}$, bokeh level $K$, and a shape kernel $s$ (binary/raster PSF), the simulator renders
$I_{\mathrm{syn}}\;=\;\mathcal{R}\!\big(I_{\mathrm{aif}},D;\,S_{1},K,s\big),$ (3)
yielding paired supervision for shape-aware training.
Shape-conditioned fine-tuning.
Starting from our base model trained without shape control, we freeze all original LoRA weights and introduce another conditioning branch that ingests a shape exemplar $s$. Only this branch is fine-tuned while the backbone remains fixed, ensuring that shape edits do not regress the learned bokeh synthesis.
4 Experiments
4.1 Setup
Backbone and training details.
Our base generative backbone is FLUX-1-dev [61], fine-tuned via LoRA [28] with a multi-condition scheme following [66]. DeblurNet employs LoRA rank r=128r{=}128, while BokehNet uses r=64r{=}64.
Both models are trained with per-GPU batch size $1$ and gradient accumulation over $8$ steps on $4{\times}$RTX A6000 GPUs. DeblurNet is trained for $60\mathrm{K}$ steps. BokehNet is trained in two stages: synthetic pair training for $40\mathrm{K}$ steps and Real unpair for another $40\mathrm{K}$ steps.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input GT BokehMe [48] Bokehlicious [60] BokehDiff [89] Ours
Figure 5: Qualitative comparison on bokeh synthesis benchmark. Results on LF-Bokeh with zoomed patches (blue and orange boxes) highlighting detail quality. Our BokehNet synthesizes bokeh effects that better match ground truth with realistic blur gradients and natural occlusion handling. Baselines show various artifacts: BokehMe [48] exhibits simulator bias, Bokehlicious [60] over-smooths details, and BokehDiff [89] produces inconsistent defocus. Our semi-supervised training on real bokeh images enables the capture of authentic lens characteristics.
Datasets.
We summarize the training data for the two modules below.
DeblurNet. We train on DPDD [1] together with a curated subset of RealBokeh_3MP [60], totaling approximately 4.5K image pairs.
BokehNet (synthetic pairs). We use AIF images from [82] and a subset of EBB! [30], and procedurally render defocus under varying aperture sizes/shapes and focus points to generate about 70K synthetic pairs.
BokehNet (real supervision). We incorporate (i) an unpaired real-bokeh set collected from Flickr by Fortes et al. [19], and (ii) real AIF–bokeh pairs with incomplete metadata from RealBokeh_3MP [60] and LFDOF [55], contributing approximately 27K real examples in total. All datasets are used strictly for research and are split so that no scene overlaps across training and evaluation.
Table 3: Bokeh synthesis benchmark. Evaluation on LF-Bokeh (200 light-field images with multiple focus planes). Following [89], we use per-image binary search for optimal $K$. Bokehlicious supports only aperture size control. Our semi-supervised training on real bokeh data outperforms all baselines.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
BokehMe [48] 0.1228 0.0744 0.9511
Bokehlicious [60] 0.1799 0.1062 0.9304
BokehDiff [89] 0.1708 0.0933 0.9192
GenRefocus (Ours) 0.1047 0.0611 0.9570
Benchmarks.
We evaluate three tasks: (i) defocus deblurring on DPDD and RealDOF; (ii) bokeh synthesis on our new LF-Bokeh benchmark, comprising $200$ images with multiple focus planes synthesized from [53] and [15] light-field captures; and (iii) refocusing on LF-Refocus, built by pairing same-scene images within LF-Bokeh to form $400$ source–target pairs.
Inference-time resolution and aspect ratio.
We adopt a tiling strategy inspired by  [5] at evaluation time: we split the input image into overlapping patches and perform tile-wise denoising before blending them back together. As a result, all benchmarks are conducted at each dataset’s original aspect ratio and resolution, without additional resizing.
4.2 Comparison
Defocus deblurring.
As stated on Tab. 2, DeblurNet consistently outperforms all baselines across reference and no-reference metrics. On certain NR-IQA measures, it even surpasses the ground-truth. Qualitatively, Fig. 4 (a) shows faithful recovery of the “NEW YORK” text with well-preserved edges. The scene in Fig. 4 (b) is particularly challenging: competing methods recover only a blurry dark stripe and further distort the background, whereas DeblurNet reconstructs structure-consistent content and delivers visually compelling results without introducing artifacts or geometry drift.
Table 4: Quantitative comparison on refocusing benchmark. Results on our LF-Refocus dataset (400 pairs with different focus planes). We compare our DeblurNet+BokehNet pipeline against combinations of existing all-in-focus estimation methods [56, 84] and bokeh synthesis methods [48, 89]. Optimal bokeh level $K$ is determined via per-image binary search following [89]. Our approach achieves superior performance across all metrics, demonstrating the benefit of joint optimization and semi-supervised training.
AiF Estimation Bokeh Synthesis LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ NIQE $\downarrow$ CLIP-IQA $\uparrow$
DRBNet [56] BokehMe [48] 0.1471 0.0937 0.9335 43.3087 6.5783 0.4460
BokehDiff [89] 0.1848 0.1030 0.9310 48.0766 5.8195 0.4997
Restormer [84] BokehMe [48] 0.1554 0.0963 0.9303 43.3188 7.1249 0.4782
BokehDiff [89] 0.1860 0.1016 0.9320 48.0326 6.3216 0.5352
DeblurNet (Ours) BokehNet (Ours) 0.1458 0.0850 0.9451 52.6954 5.3343 0.5763
Bokeh synthesis.
Following the BokehDiff protocol [89], we conduct a per-image binary search over the bokeh level $K$ and select the value that maximizes SSIM with the target. All other control signals are held fixed, ensuring a fair comparison in which each method operates at its best attainable setting. As summarized in Tab. 3, our approach attains the highest scores on both fidelity and perceptual metrics, outperforming all baselines. The qualitative results in Fig. 5 show that our method synthesizes bokeh that more closely matches that of real photographs. In the zoomed-in patches, our outputs more faithfully reproduce the ground-truth bokeh than competing baselines. Other methods tend to naively translate the predicted defocus map into blur strength, so any error in the depth scale directly leads to over- or under-blurring. In contrast, BokehNet, trained with our semi-supervised strategy, is more robust to such depth-scale inaccuracies and thus produces a more appropriate amount of blur that better matches the ground truth.
Refocusing.
We build four two-stage baselines by pairing strong AIF estimators (DRBNet [56], Restormer [84]) with bokeh synthesizers (BokehMe [48], BokehDiff [89]), yielding the $2\times 2$ combinations. For each method, we also follow the protocol of BokehDiff [89] to select the best bokeh level $K$ . As shown in Tab.4, our approach achieves the best results on both fidelity and perceptual metrics, outperforming all baseline pairings. Qualitatively, Fig. 6 shows that competing methods tend to produce spatially ambiguous blur and fail to place the focal plane correctly, resulting in uniformly soft content and inaccurate foreground/background separation. In contrast, our approach restores fine details at the intended focus while synthesizing physically consistent bokeh elsewhere.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input GT DRBNet [56] DRBNet [56] Restormer [84] Restormer [84] GenRefocus
+ BokehMe [48] + BokehDiff [89] + BokehMe [48] + BokehDiff [89] (Ours)
Figure 6: Qualitative comparison on refocusing benchmark. Results on LF-Refocus comparing our end-to-end trained pipeline against baseline combinations of existing deblurring [56, 84] and bokeh synthesis [48, 89] methods. Zoomed patches (blue and orange boxes) highlight focal and defocused regions. Baselines struggle with accurate focus-plane placement and produce inconsistent or overly smooth blur. Our method achieves sharp refocusing at the target plane while synthesizing photorealistic bokeh, closely matching ground truth across diverse scenes.
Table 5: Ablation on pipeline design. Comparing one-stage (direct refocusing) and two-stage (deblur then re-bokeh) on LF-Refocus. Both use the same backbone and synthetic data. Two-stage enables better depth control and independent semi-supervised learning per stage, yielding substantial improvements.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
One stage 0.1723 0.0976 0.9345
Two stage (ours) 0.1458 0.0850 0.9461
Table 6: Ablation on BokehNet training. Comparing synthetic-only vs. adding real unpaired bokeh images on LF-Bokeh. Second stage uses real bokeh with estimated parameters (Sec. 3.2), initialized from synthetic weights. Real data enables learning authentic optical characteristics beyond simulators.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
Synthetic pair 0.1289 0.0738 0.9461
Synthetic pair+ Real unpaired 0.1047 0.0611 0.9570
4.3 Ablation Study
One stage and two stage refocusing.
We also evaluate a variant that takes an input image and a provided defocus map and directly produces a refocused output in a single pass. Concretely, we instantiate a one-stage model by reusing the BokehNet backbone and training it on the same synthetic pairs, then assess it on the refocusing benchmark. As reported in Tab. 5, the one-stage variant lags the proposed two-stage pipeline by a substantial margin. We attribute this gap to two factors. First, control fidelity: the defocus map relies on depth estimated from a blurry input; without an all-in-focus view, depth quality degrades, leading to misaligned or biased blur fields and weakening refocusing accuracy. Second, data leverage: the one-stage formulation cannot separate the problem into defocus removal and bokeh synthesis, preventing us from injecting real-data supervision tailored to each subtask.
Key idea and BokehNet training ablation.
Our core idea is to use data generation to harvest real-world supervision for training BokehNet, leveraging real unpaired images. To quantify its impact, we evaluate two training settings without introducing additional architectural changes: (i) synthetic pair: BokehNet is trained purely on synthetic paired data; and (ii) synthetic pair + real unpaired: starting from (i), we further train with real unpaired images processed by our pipeline to produce usable supervision. On the bokeh-synthesis benchmark, the synthetic pair setting performs on par with a classical simulator baseline, while the synthetic pair + real unpaired setting delivers substantial improvements across both fidelity and perceptual metrics (Tab. 6).
Application.
As illustrated in Fig. 7, after we fine-tune BokehNet on our point-light 1K dataset, the model can synthesize bokeh with diverse, user-specified aperture shapes. Moreover, even though DeblurNet is trained without any text prompts, it can still leverage textual instructions at inference time to refine the all-in-focus output, as shown in Fig. 8. This highlights a key advantage of generative defocus-deblurring models: they naturally support post-hoc, prompt-based correction and editing of the reconstructed content.
Refer to caption Refer to caption Refer to caption Refer to caption
Input Triangle Heart Star
Figure 7: Controllable aperture shape synthesis. Given an input image, our BokehNet can synthesize bokeh effects with custom aperture shapes (triangle, heart, star) by conditioning on shape exemplars. Background point lights exhibit the specified aperture geometry while maintaining scene consistency.
Figure 8: Text-guided deblurring. DeblurNet uses optional text prompts to guide reconstruction of severely blurred text. With prompts, we correct hallucinations (red: “DESION” $\rightarrow$ “DESIGN”) and recover accurate text (green: “yeurs” $\rightarrow$ “years”), matching GT. Without guidance, it hallucinates incorrect content.
5 Conclusion
We present Generative Refocusing, a two-stage diffusion framework turning single images into controllable virtual cameras. By decoupling into DeblurNet (deblurring) and BokehNet (bokeh synthesis), our approach handles arbitrary inputs with intuitive controls over focus plane, bokeh intensity, and aperture shape. Through semi-supervised training on synthetic and real unpaired bokeh data, GenRefocus learns authentic optical behavior beyond simulators. Experiments on DPDD, RealDOF, LF-Bokeh, and LF-Refocus show consistent improvements, enabling text-guided refinement and aperture-shape editing.
Limitations and future work.
GenRefocus relies on monocular depth estimation; severe failures result in mislocalized focal planes. Complex aperture shapes require simulator-driven training data. Future work should generalize to richer aperture vocabularies, including user-drawn designs.
Appendix A Appendix Overview
In this appendix, we provide additional details on the training datasets used in our framework, the construction pipeline of our PointLight-1K dataset, a comparison with the VLM model, and visual comparisons from the BokehNet training ablation study. We also present failure cases and additional in-the-wild results on the accompanying HTML page.
Figure 9: Comparison with VLM model. Qualitative refocusing results comparing our Generative Refocusing framework with Gemini 3 Nano Banana Pro [20] given prompt ”focus on the man on the right”.
Appendix B Visual Comparison for BokehNet Training Ablation
In the main paper, we compare different training strategies for BokehNet. As illustrated in Fig. 10, under a foreground-bokeh setting with comparable bokeh strength in both rows, the model in (a), which is trained only on synthetic pairs, inherits artifacts from the simulator [48] and fails to correctly blur fine structures such as hair along object boundaries. In contrast, the model in (b), which is additionally trained with unpaired real bokeh data, produces smoother and more realistic blur around these edges.
Appendix C Training Datasets
We summarize the datasets used to train DeblurNet and BokehNet.
C.1 DeblurNet
For training DeblurNet, we use the DPDD [1] dataset and adopt all images from its official training split. This dataset provides paired defocused and all-in-focus images, which we use to supervise the deblurring stage.
We further leverage the RealBokeh_3MP [60] dataset, which contains images captured with relatively small apertures and thus does not strictly provide all-in-focus views. To obtain images that are sufficiently sharp for our purposes, we compute the Laplacian variance of each image as a simple focus measure and rank the dataset accordingly. We then retain the top $3000$ images with the highest Laplacian variance as additional training data for DeblurNet.
C.2 BokehNet
For training BokehNet, we require diverse all-in-focus inputs. We therefore draw all-in-focus images from the [82] and [30] collections. As in the DeblurNet stage, we compute the Laplacian variance for each candidate image and filter out blurry examples. After this filtering, we obtain a pool of approximately 1.71.7k all-in-focus images for simulating bokeh images. The handling of the real bokeh datasets BokehDiffusion [19] and RealBokeh_3MP [60] for bokeh synthesis is already described in the main paper.
Appendix D PointLight-1K Collection Pipeline
We construct the PointLight-1K dataset to better model scenes with strong point-light bokeh. The pipeline consists of the following steps:
  1. 1.
    Keyword mining. We first mine keywords from photographs on Flickr [^1] https://www.flickr.com/ that are tagged with terms such as “night” or “bokeh”. These keywords capture typical compositions and scene elements with prominent point lights.
  2. 2.
    Prompt expansion. The collected keywords are then expanded into diverse, natural-language prompts using GPT-4o [47]. This step enriches the textual descriptions while preserving point-light characteristics.
  3. 3.
    Image generation with fine-tuned FLUX. We use a fine-tuned version of FLUX.1-Dev [61] (FLUX.1-dev-LoRA-AntiBlur LoRA) to generate images from the prompts. The fine-tuning is designed to encourage sharper scenes.
  4. 4.
    Deblurring with DeblurNet. Despite fine-tuning, we observe that images involving point lights may still exhibit mild residual blur. To further reduce this blur, we pass all generated images through our DeblurNet module, optimizing each image toward an all-in-focus appearance.
  5. 5.
    Final dataset. After this process, we obtain a curated set of 11k night-time images with prominent point lights and minimal residual blur. We refer to this dataset as PointLight-1K and use it as a specialized training set to pretrain and refine our learning strategy for controllable bokeh shape.
Appendix E Comparison with VLM model
We also compare our method with the recently released vision–language model Gemini-3 Nano Banana Pro [20], which explicitly supports text-driven refocusing. For this model, we use the prompt “focus on the man on the right” to specify the desired focus point. As shown in Fig. 9, Gemini-3 Nano Banana Pro can change the focus to some extent, but it also alters the facial expressions and appearances of people in both the foreground and the background. In contrast, our method preserves the original content while producing a realistic bokeh effect and accurate refocusing.
Figure 10: Visual Comparison for BokehNet Training Ablation.
Appendix F Failure Cases
We observe two main types of unsatisfactory refocusing results: (1) incorrect hallucinated details and (2) inaccurate bokeh blur. For the first type, even though DeblurNet can generate visually plausible high-frequency details, in cases of severe input blur it may hallucinate incorrect detail (see Fig. 11). For the second type, the bokeh blur can be imperfect because our defocus map is derived from monocular depth estimator [7]. When the estimated depth scale is biased, the resulting defocus map may fail to fully cover the intended focus region or may assign too much or too little blur to certain areas. Although our semi-supervised training strategy for BokehNet mitigates these issues to some extent, residual discrepancies with respect to the ground truth can still be observed in challenging cases.
Figure 11: Failure case. Due to the severely blurred input, the model hallucinates an incorrect time, changing 11:30 a.m. to 12:30 a.m.

### v3

@@
 Abstract
-Depth-of-field control is essential in photography, but getting the perfect focus often takes several tries or special equipment.
+Depth-of-field control is essential in photography, but achieving perfect focus often requires multiple attempts or specialized equipment.
 Single-image refocusing is still difficult.
@@
 Current methods have significant drawbacks.
-They need all-in-focus inputs, depend on synthetic data from simulators, and have limited control over aperture.
-We introduce Generative Refocusing, a two-step process that uses DeblurNet to recover all-in-focus images from various inputs and BokehNet for creating controllable bokeh.
-Our main innovation is semi-supervised training.
-This method combines synthetic paired data with unpaired real bokeh images, using EXIF metadata to capture real optical characteristics beyond what simulators can provide.
+They require all-in-focus inputs, rely on synthetic data from simulators, and have limited control over the aperture.
+We introduce Generative Refocusing, a two-step process that uses DeblurNet to recover all-in-focus images from diverse inputs and BokehNet to create controllable bokeh.
+This method combines synthetic and real bokeh images to achieve precise control while preserving authentic optical characteristics.
 Our experiments show we achieve top performance in defocus deblurring, bokeh synthesis, and refocusing benchmarks.
-Additionally, our Generative Refocusing allows text-guided adjustments and custom aperture shapes.
+Additionally, our Generative Refocusing allows custom aperture shapes.
 Project page: https://generative-refocusing.github.io/
+Figure 1: Generative refocusing with controllable bokeh.
+Our model turns a single input image into a virtual camera that users can control.
+This enables a range of adjustments after image capture.
+(a) demonstrates aperture size control, allowing the user to change the depth of field from strong bokeh to an all-in-focus image.
+(b) demonstrates focus plane control by shifting the sharp region from the middle subject to the background.
+(c) highlights aperture-shape control, synthesizing a creative heart-shaped bokeh from point lights in the scene.
+(d) shows composite control, where both the focus plane and aperture size are adjusted simultaneously to reframe the subject.
 1 Introduction
@@
 It helps guide the viewer’s eye through selective focus and bokeh effects.
-Getting perfect focus often requires multiple shots or special equipment, which can be tough for everyday photographers.
+Achieving perfect focus often requires multiple shots or specialized equipment, which can be challenging for everyday photographers.
 Single-image refocusing removes these obstacles.
@@
 The main challenge is recovering sharp details from blurred areas while creating natural-looking bokeh, while still giving users fine control.
-Existing research addresses different aspects of this problem.
+Existing research has addressed various aspects of this problem.
 Defocus deblurring methods [1, 8, 32, 53, 65, 71, 82] aim to recover sharp images.
@@
 End-to-end refocusing often requires specialized capture setups [35, 41, 2].
-Despite these advances, three main limitations remain (Tab.
-1).
+Despite these advances, three main limitations remain (Tab.˜1).
 First, most methods rely on all-in-focus inputs and accurate depth maps.
 This restricts their use to images that already have defocus blur.
-Second, while synthetic data [48, 89] can be used for training, it is limited by the realism of the simulators.
+Second, while synthetic data [43, 87, 69] can be used for training, its realism is limited by the fidelity of the simulators.
 This approach often fails to capture detailed visual appearances.
-On the other hand, obtaining real paired datasets [60] is very expensive.
-Third, current methods usually only support aperture size, not its shape.
+On the other hand, real datasets [57, 64] suffer from a critical limitation: they are restricted to isolated one-dimensional variations (i.e., solely aperture sweeps or focus plane sweeps).
+Consequently, they fail to provide the joint distribution of focus and aperture changes required to train a model for flexible, dual-parameter refocusing.
+Third, current methods typically support only aperture size, not its shape.
 This restriction limits the possibilities for creative bokeh effects.
 Table 1: Comparison across methods.
-= supported;  = not $\LARGE\triangle = supported$ with caveat.
-In the synthetic/real training data columns, “paired” indicates the use of paired supervision, while “unpaired” indicates real unpaired data (no real paired supervision).
+= supported;  = not supported; $\triangle$} = supported with caveat.
+Flexible input indicates compatibility with both all-in-focus (AIF) and defocused images.
+BokehMe and Bokehdiff require AIF inputs, while Learn2refocus strictly uses defocused inputs.
+For BokehMe, only the classical renderer variant supports aperture-shape control.
 Bokehlicious typically assumes an all-in-focus input and requires fine-tuning for different input types.
-For BokehMe, only the classical renderer variant supports aperture-shape control.
-UW = ultra-wide, W = wide.
-Method Input Training data Supported user control
-Setup AIF required? Synthetic Real Aperture size Aperture shape Focus point Defocus map
-[…]
-DC2 [2] Dual-view (UW+W) paired
-Bokehlicious [60] Single-view paired
-BokehDiff [89] Single-view paired
-DiffCamera [71] Single-view paired
-Ours Single-view paired unpaired
-To tackle these issues, we present Generative Refocusing, a flexible single-image refocusing system built on a two-stage design.
+Learn2Refocus generates focal stacks at discrete depth planes rather than supporting point-based focus selection.
+Method Flexible [1pt]input? Synthetic training data Real training data Supported user control types
+Focus- [1pt]sweep Aperture- [-1pt]sweep Single [-3pt]images Aperture [-3pt]size Aperture [-3pt]shape Focus [-1pt]point Defocus map
+BokehMe [43] $\triangle$}
+Bokehlicious [57]
+BokehDiff [87]
+DiffCamera [69]
+Learn2Refocus [64] $\triangle$}
+GenRefocus (Ours)
+To tackle these issues, we present Generative Refocusing (GenRefocus), a flexible single-image refocusing system built on a two-stage design.
 Our DeblurNet module produces a sharp, in-focus image from a variety of inputs.
 It uses a diffusion model guided by initial deblurring predictions.
-Our BokehNet creates fully customizable bokeh, considering user-defined focus planes, bokeh intensity, and aperture shapes (LABEL:fig:teaser).
-A significant advancement in our method is our training scheme, which makes use of both 1) paired synthetic data to maintain geometric consistency and 2) unpaired real-world bokeh images with camera metadata that help us learn real lens characteristics not captured by simulators.
+Our BokehNet generates fully customizable bokeh, accounting for user-defined focus planes, bokeh intensity, and aperture shapes (Fig.˜1).
+A significant advancement of our method is the proposed training scheme.
+While we utilize synthetic data to maintain geometric consistency, our key contribution lies in handling real-world data.
+By leveraging auxiliary information and dedicated processing, we complete the missing input or control signals (i.e., defocus maps) in existing real-world datasets.
+This strategy enables our model to effectively learn authentic lens characteristics that simulators fail to capture.
 Our approach achieves top performance in defocus deblurring, bokeh generation, and refocusing, while maintaining the natural consistency of scenes.
 Our main contributions are:
-  • A flexible refocusing pipeline accepts images in any focus state. It gives users control over the focus plane, bokeh intensity, and aperture shape through two-stage decomposition.
-  • This semi-supervised framework connects synthetic and unpaired real bokeh images.
-It enables genuine optical learning that goes beyond what a simulator can offer.
-This is achieved through parameter estimation from EXIF metadata.
+  • A flexible refocusing pipeline accepts images in any focus state. It provides users with control over the focus plane, bokeh intensity, and aperture shape via two-stage decomposition.
+  • A novel training scheme that overcomes the limitations of existing real-world datasets by completing absent input or control signals.
+This strategy successfully unites the geometric consistency of synthetic data with the authentic aberrations of real lenses.
   • The system performs well across defocus deblurring, bokeh synthesis, and refocusing.
 This has been validated on existing benchmarks (DPDD [1], RealDOF [32]) and our new light-field datasets (LF-Bokeh, LF-Refocus).
-It also has applications in text-guided deblurring and creating aperture shapes.
+It also has applications in creating aperture shapes.
 2 Related Work
@@
 Our method decomposes single-image refocusing into two stages: (a) Defocus Deblurring and (b) Shallow Depth-of-Field Synthesis.
-Given a blurry input image $I_{\text{in}}$, we optionally apply a pre-deblurring method [56] to obtain a conservative estimate $I_{\text{syn}}$, then feed both $I_{\text{in}}$ and $I_{\text{syn}}$ into DeblurNet to recover a high-quality all-in-focus image $I_{\text{aif}}$.
-The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ convert images to latent representations for processing by the DiT backbone.
-In the second stage, BokehNet takes the all-in-focus image $I_{\text{aif}}$, the defocus map $D_{\text{def}}$, and optionally a specific aperture shape as inputs to synthesize the refocused output.
-The defocus map $D_{\text{def}}$ is computed based on the estimated depth map $D$ [7], along with both the user-specified focus plane $S_{1}$ and bokeh level $K$.
-Diffusion Models for Image Restoration.
+(a) The noisy latent $X_{t}$ and the encoded blurry input $I_{\text{in}}$ are concatenated into a unified token sequence $\mathbf{S}_{t}$.
+DeblurNet iteratively denoises this sequence to reconstruct a high-quality all-in-focus image $I_{\text{aif}}$.
+(b) Following the same unified-sequence design, BokehNet conditions on $I_{\text{aif}}$ and the defocus map $D_{\text{def}}$ to synthesize the refocused output $I_{\text{out}}$, with an optional aperture-shape condition AA.
+The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ map images to latent representations for the DiT backbone.
+The defocus map $D_{\text{def}}$ is computed from the estimated depth map $D$ [7] together with the user-specified focus plane $D_{\text{focus}}$ and bokeh level $K$.
+2.0.1 Diffusion Models for Image Restoration.
 Diffusion models [22, 61, 79, 66] advance image restoration via generative priors, extending to zero-shot restoration and artifact removal.
@@
 ResShift [81] cuts steps from 1000 to 4–15 via residual shifting.
-Deblurring progressed from merging diffusion priors with regression tasks, such as matting [74] or hierarchical restoration [13], to learning spatially-varying kernels in latent space [33].
+Deblurring has progressed from merging diffusion priors with regression tasks, such as matting [72] or hierarchical restoration [12], to learning spatially varying kernels in latent space [30].
 Cascaded pipelines [23] refine at multiple resolutions, while principled frameworks [40, 39] treat degradations as stochastic processes.
@@
 Unlike these general methods, we target spatially-varying defocus blur with a two-stage pipeline explicitly separating deblurring from bokeh synthesis for controllable refocusing.
-Defocus Deblurring.
+2.0.2 Defocus Deblurring.
 Defocus deblurring recovers sharp images from spatially varying blur caused by limited depth of field.
@@
 While these excel as standalone solutions, we position deblurring as the first stage of refocusing, using FLUX’s [58] generative prior to enable controllable bokeh synthesis from any input.
-Bokeh Rendering.
+2.0.3 Bokeh Rendering.
 Bokeh rendering progressed from physically based scattering [45, 34, 33] and differentiable rendering [60] to neural methods combining ray tracing with learned enhancements [43, 44].
-Learning-based approaches evolved from fixed apertures [30] to variable f-stops [60, 9], extending to video [80] and 3D scenes [62, 72].
+Learning-based approaches evolved from fixed apertures [27] to variable f-stops [57, 9], extending to video [78] and 3D scenes [59, 70, 26].
 Diffusion models enabled bokeh control [19, 80], mainly for text-to-image synthesis.
 These methods require all-in-focus inputs and accurate depth.
-We remove this constraint via flexible input handling, which accepts defocused or all-in-focus images, and semi-supervised learning on unpaired real-world bokeh to capture authentic lens characteristics beyond simulators.
-Single-Image Refocusing.
-Post-capture refocusing manipulates focus after capture.
-Light-field cameras [46, 40] enable optical refocusing but need specialized capture.
-Computational approaches evolved from deconvolution [4, 86] producing artifacts, to GAN-based paradigms [58], to diffusion methods [71].
-3D representations [62, 72, 35, 43] achieve real-time controllable DoF but need multi-view capture or calibration.
-We combine three capabilities: (1) semi-supervised learning mixing synthetic and unpaired real data for authentic optics, (2) single-image flexibility accepting any input without preprocessing, and (3) comprehensive control over focus, intensity, and aperture shape.
-Camera-Conditioned Diffusion.
+We relax this constraint via flexible input handling, which accepts defocused or fully in-focus images.
+2.0.4 Single-Image Refocusing.
+Post-capture refocusing evolved from deconvolution [4, 84] and GANs [55] to recent diffusion models [46], yet significant limitations remain (see Tab.˜1).
+Constrained by conditioning on explicit focus coordinates, DiffCamera [69] requires fixed-resolution inputs ($512\times 512$) and suffers performance drops on arbitrary aspect ratios.
+Furthermore, it inherits simulator artifacts from its purely synthetic training.
+Learn2Refocus [64] trains on real data but lacks aperture-size control and generates a focal stack at predefined depth planes that may not align with the user’s target focus.
+Additionally, 3D representations [59, 70, 31, 38] offer controllable DoF but require multi-view capture.
+To address these gaps, we combine three capabilities: (1) Hybrid training scheme mixing synthetic and real data for authentic optics, (2) single-image flexibility accepting any input without preprocessing, and (3) comprehensive control over focus, intensity, and aperture shape.
+2.0.5 Camera-Conditioned Diffusion.
 Conditioning diffusion models on camera parameters progressed from extrinsic control for view synthesis [24, 13] and video generation [21, 73], to intrinsic control of focal length, aperture, and exposure [18, 11].
@@
 Recent work uses temporal modeling [80], explicit calibration [6, 15], or transformer-based pose handling [3, 56].
-These focus on generation; we operate in editing, manipulating existing images with scene consistency.
-By decoupling into deblurring and bokeh stages, we achieve efficient single-image manipulation without multi-frame overhead.
-Semi-Supervised Learning.
-When paired data is scarce, semi-supervised learning provides alternatives.
-The field evolved from learning without clean targets [39, 34, 27] to pseudo-labeling [11, 23, 69].
-Foundation models enabled the generation of high-quality pseudo-labels [21, 85].
-Unpaired frameworks [31, 90, 64] learn enhancement without paired supervision.
-Two-stage approaches [29, 78] bridge fidelity and perceptual quality via different supervision types.
-We adapt these to bokeh rendering: perfect paired refocusing data is nearly impossible to obtain.
-We train on synthetic pairs for structure, then unpaired real bokeh for realism.
-Deblur restores to all-in-focus, enabling bokeh to learn authentic lens characteristics exceeding simulators.
+These focus on generation; we operate in editing, manipulating images with scene consistency.
+By decoupling deblurring and bokeh stages, we achieve efficient single-image manipulation without multi-frame overhead.
 3 Method
@@
 When the input is blurry, monocular depth estimation is brittle, undermining precise control over the defocus map and often yielding misfocus or artifacts.
-We propose Generative Refocusing, which decomposes the task into defocus deblurring and bokeh synthesis, and comprises two models, DeblurNet and BokehNet, that jointly enable precise and controllable refocusing (see Fig.
+We propose Generative Refocusing (GenRefocus), which decomposes the task into defocus deblurring and bokeh synthesis, and comprises two models, DeblurNet and BokehNet, that jointly enable precise and controllable refocusing (see Fig.
 2 for an overview of the pipeline); Such a flexible pipeline allows us to exert greater control over input conditions (see Tab.
@@
 Figure 3: Training data generation.
-Each training sample consists of the following five components: (i) a bokeh image, (ii) an all-in-focus (AIF) image, (iii) a depth map $D$, (iv) a bokeh level $K$, and (v) a focus plane $S_{1}$.
-We construct these samples via three routes: (a) Synthetic paired data.
-Given real AIF images and depth maps $D$, we compute a defocus map $D_{\text{def}}$ parameterized by a specified bokeh level $K$ and focus plane $S_{1}$, and feed it into a bokeh renderer [48] to synthesize corresponding bokeh images.
-(b) Real unpaired data.
+Each training sample consists of five components: (i) a bokeh image, (ii) an all-in-focus (AIF) image, (iii) a depth map $D$, (iv) a bokeh level $K$, and (v) a focus plane $D_{\text{focus}}$.
+We construct these samples via three routes: (a) Synthetic data.
+Given real AIF images and depth maps $D$, we compute a defocus map $D_{\text{def}}$ parameterized by randomly selected bokeh level $K$ and focus plane $D_{\text{focus}}$, and feed it into a bokeh renderer [43] to synthesize corresponding bokeh images.
+(b) ITW dataset [19].
 Given real bokeh images, DeblurNet recovers an AIF image.
-We then estimate depth and extract a foreground mask [88] to define the focus plane $S_{1}$.
-The bokeh level $K$ is computed from the EXIF metadata following the formulation in [19].
-(c) Real paired data without EXIF.
-For real pairs lacking EXIF, we obtain $S_{1}$ as in (b), and follow Eq.
+We then estimate depth and extract a foreground mask [86] to define the estimated focus plane $D_{\text{focus}}$.
+The bokeh level $K$ is computed from the EXIF metadata and the estimated $D_{\text{focus}}$ following the formulation in [19].
+(c) LFDOF [52] and RealBokeh.
+For real pairs, we obtain $D_{\text{focus}}$ as in (b), and follow Eq.
 (2) to estimate the bokeh level $K$.
 3.1 Stage 1: Defocus Deblurring
-DeblurNet.
-We condition the deblurring network on a pair $(I_{\text{in}},I_{\text{pd}})$, where $I_{\text{pd}}$ is a conservative pre-deblur generated by an off-the-shelf defocus deblurring model.
-In our implementation, we use a non-generative defocus model  [56], whose outputs are typically content-faithful yet over-smoothed.
-Our model, guided by a learned diffusion prior, then injects high-fidelity details without sacrificing geometric correctness.
-Artifact-robust dual conditioning.
-To prevent artifacts in $I_{\text{pd}}$ from biasing the reconstruction and to ensure that fine details are recovered from $I_{\text{in}}$, we introduce two complementary mechanisms (see Fig. 2 (a)):
-  • Positional disentanglement.
-We encode the two inputs with disjoint 2D positional grids: $I_{\mathrm{in}}$ is anchored at $(0,0)$, while $I_{\mathrm{pd}}$ is shifted to start at $(0,N)$, where $N$ denotes the side length of the token grid after downsampling $I_{\mathrm{in}}$.
-This shift lets the network distinguish the two sources, using $I_{\mathrm{pd}}$ as a low-frequency anchor while extracting high-frequency evidence from $I_{\mathrm{in}}$.
-  • Pre-deblur dropout.
-During training, we stochastically replace $I_{\mathrm{pd}}$ with a zero tensor (at a fixed probability), forcing the network to remain effective when the auxiliary input is unreliable or absent.
-This regularizes the fusion, shields against method-specific artifacts, and enables the model to operate in a single-input deblurring mode at test time.
+Given a blurry input image $I_{\text{in}}$, we follow the conditioning design of [62, 63]. At each denoising step $t$, DeblurNet processes a single unified token sequence
+$\small\mathbf{S}_{t}=\bigl[\,X_{t}\,;\,\mathcal{E}(I_{\text{in}})\,\bigr],$ (1)
+where $X_{t}$ denotes the noisy image latent tokens at timestep $t$, $\mathcal{E}$ is a pretrained VAE encoder that maps the input image to latent condition tokens, and $\cdot\,$$\cdot\,$] denotes token concatenation.
+After iterative denoising, we decode the image latent part with the VAE decoder $\mathcal{D}$ to obtain the all-in-focus output $I_{\text{aif}}$ (Fig.
+2 (a)).
+We train DeblurNet in a supervised manner using real paired data [1, 57].
 3.2 Stage 2: Shallow Depth-of-Field Synthesis
-Synthetic-pair learning.
-BokehNet expects an AIF image, a target bokeh image, and the corresponding defocus map (see Fig.
+Given an all-in-focus image $I_{\text{aif}}$ and a user-specified focus plane $D_{\text{focus}}$ with bokeh level $K$, we define the defocus map $D_{\text{def}}$ as
+$\small D_{\text{def}}=K\cdot\left|D-D_{\text{focus}}\right|,$ (2)
+where $D$ is the monocular depth map estimated from $I_{\text{aif}}$ using an off-the-shelf depth estimator [7].
+Following the unified-token conditioning design in Stage 1, BokehNet concatenates the condition tokens of $I_{\text{aif}}$ and $D_{\text{def}}$ with the noisy latent tokens at each denoising step.
+BokehNet then iteratively denoises $X_{t}$ to synthesize the refocused output $I_{\text{out}}$, as illustrated in Fig.
+2 (b).
+3.2.1 Training data and supervision.
+Training BokehNet requires $(I_{\text{aif}},I_{\text{out}},D_{\text{def}})$.
+Because accurately paired real supervision is scarce, we construct supervision via three complementary sources (Fig.
+3): (a) synthetic data, (b) ITW dataset[19], and (c) LFDOF[52] and RealBokeh[57].
+(a) Synthetic data.
+We pretrain with synthetic data: starting from real all-in-focus images and their estimated depth map $D$, we randomly sample a focus plane $D_{\text{focus}}$ and a target bokeh level $K$ (Fig.
+3 (a)).
+We then construct $D_{\text{def}}$ using Eq.
+2 and use a simulator [43] to render the corresponding target bokeh image consistent with $D_{\text{def}}$.
+This synthetic pretraining helps the network modulate the circle of confusion according to $D_{\text{def}}$; however, it is constrained by renderer bias and may introduce unrealistic artifacts.
+(b) ITW dataset.
+To capture real optics, we leverage ITW dataset[19] contains real bokeh images (see Fig.
 3 (b)).
-In practice, perfectly paired data and accurate defocus maps are hard to acquire.
-We therefore pretrain with synthetic pairs: starting from real all-in-focus images and their estimated depth, we randomly sample a focus plane $S_{1}$ and a target bokeh level $K$, and use a simulator [48] to render bokeh consistent with the induced defocus map (see Fig.
-3 (a)).
-This stage quickly trains the network to modulate the circle of confusion based on the defocus map, but it remains constrained by renderer bias and may introduce non-realistic artifacts.
-Real unpaired bokeh learning. To capture real optics, we further leverage unpaired real bokeh images with EXIF metadata (see Fig. 3 (b)) and estimate an approximate bokeh level K\!K following [19]:
-$K\;\approx\;\frac{f^{2}\,S_{1}}{\,2\,F\,(S_{1}-f)\,}\times\textit{pixel\_ratio},$ (1)
-where $f$ is the focal length and $F$ denotes the aperture $f$-number; and pixel_ratio compensates for sensor size and image resolution differences.
-The AIF input is provided by DeblurNet.
-Although $S_{1}$ can sometimes be read from metadata, both estimated and true depths may deviate in real scenes.
-To obtain more reliable supervision, we estimate in-focus masks using BiRefNet [88], and have human experts filter erroneous cases and complete partial masks.
-These masks refine the focus plane and yield a consistent defocus map, enabling BokehNet to learn from real bokeh images.
-Real paired bokeh images without metadata.
-Previous bokeh-synthesis datasets [60] [30] provide real pairs but omit EXIF metadata or provide insufficient fields to estimate a reliable $K$ (see Fig.
+For each real image, we use DeblurNet to produce the AIF input and compute an approximate bokeh level K\!K following [19]:
+$\small K\;\approx\;\frac{f^{2}\,D_{\text{focus}}}{\,2\,F\,(D_{\text{focus}}-f)\,}\times\text{pixel\_ratio},$ (3)
+where $f$ is the focal length and $F$ is the aperture $f$-number, both directly obtained from EXIF metadata.
+The pixel_ratio term accounts for differences in camera sensor size and image resolution.
+Although some devices may provide a focus-distance field in EXIF, it is frequently missing or noisy; therefore, we do not rely on EXIF for $D_{\text{focus}}$.
+Instead, we compute the focus plane based on an in-focus mask and a monocular depth estimate, following the approach of  [19].
+Specifically, we estimate an in-focus mask $M$ using BiRefNet [86] and produce a depth map $D$ using a monocular depth estimator.
+The focus plane is then determined as the median depth within the masked in-focus area:
+$\small D_{\text{focus}}=\mathrm{median}\bigl(D[M]\bigr).$ (4)
+We then generate $D_{\text{def}}$ using Eq.
+2, enabling BokehNet to effectively train on real single images.
+While these estimated AIF images and defocus maps contain inevitable noise, training on them inherently aligns with our single-image refocusing pipeline, making the model more robust to depth and deblurring artifacts.
+(c) LFDOF and RealBokeh.
+These datasets provide pairs but omit EXIF metadata or provide insufficient fields to estimate $K$ (see Fig.
 3 (c)).
-We therefore adopt a simulator-in-the-loop calibration of the bokeh level: given an AIF image $I_{\mathrm{aif}}$, an estimated depth $D$, and a focus annotation obtained as in the unpaired setting (BiRefNet-based in-focus masks with expert filtering), we sweep $K$ and choose the value whose rendered result best matches the real bokeh target $I_{\mathrm{real}}$:
-$K^{\star}=\operatorname*{argmax}_{\begin{subarray}{c}K\in\\ (K_{\min},\,K_{\max})\end{subarray}}\operatorname{SSIM}\!\big(B(I_{\mathrm{aif}},D;\,S_{1},K),\,I_{\mathrm{real}}\big),$ (2)
-where B denotes our physically guided renderer and $S_{1}$ is the focus-plane proxy.
-The selected $K^{\star}$ is then used as the pseudo bokeh level label for training .
-Focus-plane annotation follows the same protocol as in the unpaired case.
-Discussion.
-Although the resulting defocus maps and the AIF proxy are not perfectly accurate, this supervision regime mirrors the practical refocusing scenario.
-Aggregating numbers of real bokeh images substantially improves realism and robustness over synthetic-only training.
+To address this, we first compute the focus-plane proxy $D_{\text{focus}}$.
+Similar to (b), we employ BiRefNet [86] to obtain an initial in-focus mask $M$.
+However, due to the increased diversity and complexity of the scenes in these datasets, the initial estimate of $M$ is sometimes unreliable.
+Rather than simply verifying and discarding unreliable cases, we introduce a manual refinement step.
+Specifically, we re-select a small yet reliable in-focus region to correct $M$, thereby accurately extracting $D_{\text{focus}}$.
+This strategy preserves challenging samples rather than excluding them, thereby enabling the model to learn from a wider range of real-world scenarios.
+With $D_{\text{focus}}$ established, we adopt a simulator-in-the-loop calibration of the bokeh level.
+Given an AIF image $I_{\mathrm{aif}}$ and estimated depth $D$, we sweep $K$ and choose the value whose rendered result best matches the real bokeh target $I_{\mathrm{real}}$:
+$\small K^{\star}=\operatorname*{argmax}_{\begin{subarray}{c}K\in\\ (K_{\min},\,K_{\max})\end{subarray}}\operatorname{SSIM}\!\big(R(I_{\mathrm{aif}},D;\,D_{\text{focus}},K),\,I_{\mathrm{real}}\big),$ (5)
+where $R$ denotes our physically guided renderer.
+The selected $K^{\star}$ is then used as the pseudo-bokeh-level label for training, provided that its corresponding SSIM exceeds a predefined threshold to ensure reliable supervision.
+Finally, the defocus map $D_{\text{def}}$ is constructed accordingly following Eq.
+2.
 3.3 Bokeh-Shape Aware Synthesis
@@
 Table 2: Defocus deblurring benchmark on RealDOF [32] and DPDD [1] datasets.
-We evaluate the task of recovering all-in-focus images from defocused-blur inputs.
-Our method, GenRefocus, uses the DeblurNet module for this task.
-We report both reference-based metrics (LPIPS, FID) and no-reference perceptual quality metrics (CLIP-IQA, MANIQA, MUSIQ).
+We report reference (LPIPS, DISTS) and no-reference (CLIP-IQA, MANIQA, MUSIQ) perceptual metrics.
 Best, second best, and third best results are highlighted.
-GenRefocus outperforms all competing approaches across both benchmarks and all metrics, demonstrating the effectiveness of our diffusion-based deblurring with pre-deblur conditioning.
 RealDOF [32] DPDD [1]
-Method LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
-Input 0.5241 92.93 0.3562 0.2213 28.7087 0.3485 88.85 0.4337 0.3325 45.5376
-GT 0.4233 0.2675 42.5777 0.4563 0.3294 47.8102
-AIFNet [55] 0.4132 72.38 0.3814 0.2536 35.7121 0.3090 79.84 0.4359 0.3371 46.9558
-IFANet [36] 0.2994 33.21 0.4088 0.2609 39.5949 0.2301 42.60 0.4472 0.3276 47.2784
-DRBNet [56] 0.2550 29.89 0.3889 0.2609 38.9014 0.1819 39.37 0.4142 0.3254 46.7229
-Restormer [84] 0.2863 29.80 0.4062 0.2642 40.8355 0.1762 33.45 0.4332 0.3307 47.9872
-INIKNet [51] 0.2851 32.18 0.3984 0.2580 38.9959 0.1860 36.63 0.4290 0.3265 47.1961
-GenRefocus (Ours) 0.2356 24.73 0.4575 0.2886 43.7042 0.1598 33.08 0.4619 0.3411 49.3019
+Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
+Input 0.5241 0.2865 0.3562 0.2213 28.7087 0.3485 0.1827 0.4337 0.3325 45.5376
+[…]
+[…]
+INIKNet [47] 0.2851 0.1615 0.3984 0.2580 38.9959 0.1860 0.1235 0.4290 0.3265 47.1961
+[…]
+[…]
+[…]
 Figure 4: Qualitative comparison on defocus deblurring.
@@
 Blue boxes on the left indicate cropped regions shown in detail.
-Our DeblurNet faithfully recovers fine text details (“NEW YORK”) in (a) where competing methods produce blurry or distorted results.
-In the challenging example (b) with severe defocus blur, other methods either fail to recover structure (AIFNet [55], DRBNet [56], INIKNet [51]) or introduce artifacts in the background (IFANet [36], Restormer [84]), while our method reconstructs geometrically consistent, visually compelling content by leveraging diffusion priors guided by the pre-deblurred input.
-Simulator and Data.
+In (a), most methods fail to resolve the text, whereas both our method and Bokehlicious [57] successfully recover the words “WELCOME” and “OPEN HOURS”.
+However, Bokehlicious produces noticeable distortions in the right-side characters (e.g., “ME” and “RS”), whereas our approach faithfully restores the original text structure.
+For the challenging example (b) with severe defocus blur, all baseline methods remain highly blurry and fail to recover meaningful details.
+In contrast, our method synthesizes a relatively clear and visually compelling output.
+3.3.1 Simulator and Data.
 Real photographs exhibiting diverse bokeh shapes are rare, and paired AIF–bokeh examples are even scarcer.
@@
 A key observation is that when an all-in-focus (AIF) image lacks point-light stimuli, the simulated bokeh carries weak shape cues, making the model reluctant to learn shape-conditioned responses.
-To address this, we synthesize a PointLight-1K AIF set designed to reveal aperture shape.
-We also extend the classical BokehMe renderer to scatter through a given shape: given an AIF image $I_{\mathrm{aif}}$, depth $D$, focus-plane proxy $S_{1}$, bokeh level $K$, and a shape kernel $s$ (binary/raster PSF), the simulator renders
-$I_{\mathrm{syn}}\;=\;\mathcal{R}\!\big(I_{\mathrm{aif}},D;\,S_{1},K,s\big),$ (3)
+To address this, we synthesize a PointLight-1K dataset designed to reveal aperture shape.
+We also extend the classical BokehMe renderer to scatter through a given shape: given an AIF image $I_{\mathrm{aif}}$, depth $D$, focus-plane proxy $D_{\text{focus}}$, bokeh level $K$, and a shape kernel $s$ (binary/raster PSF), the simulator R renders
+$\small I_{\mathrm{syn}}\;=\;\mathcal{R}\!\big(I_{\mathrm{aif}},D;\,D_{\text{focus}},K,s\big),$ (6)
 yielding paired supervision for shape-aware training.
-Shape-conditioned fine-tuning.
-Starting from our base model trained without shape control, we freeze all original LoRA weights and introduce another conditioning branch that ingests a shape exemplar $s$.
-Only this branch is fine-tuned while the backbone remains fixed, ensuring that shape edits do not regress the learned bokeh synthesis.
+3.3.2 Shape-conditioned fine-tuning.
+To incorporate the aperture shape $A$, we append its tokens directly to the unified sequence.
+To ensure that shape edits do not regress the learned bokeh synthesis, we freeze all original LoRA weights and introduce a new, trainable LoRA module.
+Only this dedicated LoRA is fine-tuned to handle the shape conditioning.
+3.4 Pre-deblur Module
+We also study a variant that augments DeblurNet with a conservative pre-deblur prior.
+We train this variant using the same data as the base DeblurNet.
+Specifically, it condition on $I_{\text{in}}$ and $I_{\text{pd}}$, where $I_{\text{pd}}$ is generated by an off-the-shelf defocus deblurring model.
+In the implementation, we use a non-generative defocus model [53], whose outputs are typically content-faithful but overly smooth.
+Qualitative results for this variant are shown in the supplementary material.
+3.5 Implementation Details
+While our backbone [58] intrinsically supports arbitrary resolutions and aspect ratios without resizing, and our conditioning setup is completely coordinate-independent, ultra-high-resolution inputs (e.g., 4K) still incur prohibitive VRAM consumption and quality degradation.
+We implement a tiling strategy inspired by [5] during inference.
+Specifically, we process the image as overlapping patches and blend the tile-wise denoised results.
+This strategy enables us to process images exactly at their original resolutions, fully preserving image fidelity.
 4 Experiments
 4.1 Setup
-Backbone and training details.
-Our base generative backbone is FLUX-1-dev [61], fine-tuned via LoRA [28] with a multi-condition scheme following [66]. DeblurNet employs LoRA rank r=128r{=}128, while BokehNet uses r=64r{=}64.
-Both models are trained with per-GPU batch size $1$ and gradient accumulation over $8$ steps on $4{\times}$RTX A6000 GPUs.
+4.1.1 Backbone and training details.
+Our backbone is FLUX-1-dev [58], fine-tuned via LoRA [25] with a conditioning scheme following [62, 63].
+DeblurNet employs LoRA rank r=128r{=}128, while BokehNet uses r=64r{=}64.
+Both models are trained with a per-GPU batch size of $1$ and gradient accumulation over $8$ steps on $4{\times}$ RTX A6000 GPUs.
 DeblurNet is trained for $60\mathrm{K}$ steps.
-BokehNet is trained in two stages: synthetic pair training for $40\mathrm{K}$ steps and Real unpair for another $40\mathrm{K}$ steps.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input GT BokehMe [48] Bokehlicious [60] BokehDiff [89] Ours
+BokehNet is trained in two stages: (i) $40\mathrm{K}$ steps on synthetic data, and (ii) $60\mathrm{K}$ steps on real data.
+4.1.2 Runtime and computational cost.
+During inference, each stage employs 28 denoising steps.
+Without the tiling strategy, our pipeline consumes approximately 40 GB of VRAM and takes approximately 50 seconds per image on a single NVIDIA RTX A6000 GPU.
+4.1.3 Datasets.
+We train DeblurNet on 3.5K pairs sourced from DPDD [1] and a subset of RealBokeh [57].
+For BokehNet, to simultaneously maintain realism and controllability, we utilize a hybrid dataset comprising $\sim$70K synthetic pairs derived from [80, 27], combined with approximately 26K real examples sourced from ITW dataset [19], RealBokeh [57], and LFDOF [52].
+Table 3: Bokeh synthesis benchmark.
+Evaluation on LF-Bokeh.
+Following [87, 43], we use per-image binary search for optimal $K$.
+Since Bokehlicious [57] uses f-stop rather than $K$ as its control input, we perform a binary search over f-stop values.
+Due to this fundamentally different control mechanism, we exclude it from the LVCorr evaluation.
+Note that Bokeh diffusion [19] targets the text-to-image task.
+To adapt it for image editing, we use textual inversion [61].
+Fidelity Controllability
+Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ LVCorr $\uparrow$
+[…]
+Bokehlicious [57] 0.1799 0.1062 0.9304 -
+[…]
+Bokeh Diffusion [19] 0.4999 0.2529 0.7569 0.8954
+[…]
+[…]
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Input GT BokehMe Bokehlicious BokehDiff Diffcamera GenRefocus
+[43] [57] [87] [69] (Ours)
 Figure 5: Qualitative comparison on bokeh synthesis benchmark.
 Results on LF-Bokeh with zoomed patches (blue and orange boxes) highlighting detail quality.
-Our BokehNet synthesizes bokeh effects that better match ground truth with realistic blur gradients and natural occlusion handling.
-Baselines show various artifacts: BokehMe [48] exhibits simulator bias, Bokehlicious [60] over-smooths details, and BokehDiff [89] produces inconsistent defocus.
-Our semi-supervised training on real bokeh images enables the capture of authentic lens characteristics.
-Datasets.
-We summarize the training data for the two modules below.
-DeblurNet. We train on DPDD [1] together with a curated subset of RealBokeh_3MP [60], totaling approximately 4.5K image pairs.
-BokehNet (synthetic pairs).
-We use AIF images from [82] and a subset of EBB!
-[30], and procedurally render defocus under varying aperture sizes/shapes and focus points to generate about 70K synthetic pairs.
-BokehNet (real supervision).
-We incorporate (i) an unpaired real-bokeh set collected from Flickr by Fortes et al.
-[19], and (ii) real AIF–bokeh pairs with incomplete metadata from RealBokeh_3MP [60] and LFDOF [55], contributing approximately 27K real examples in total.
-All datasets are used strictly for research and are split so that no scene overlaps across training and evaluation.
-Table 3: Bokeh synthesis benchmark.
-Evaluation on LF-Bokeh (200 light-field images with multiple focus planes).
-Following [89], we use per-image binary search for optimal $K$.
-Bokehlicious supports only aperture size control.
-Our semi-supervised training on real bokeh data outperforms all baselines.
+Our method synthesizes bokeh effects that better match ground truth with realistic blur gradients and natural occlusion handling.
+4.1.4 Benchmarks and Metrics
+We evaluate our method on three tasks using reference-based fidelity and no-reference image quality assessment metrics.
+(i) Defocus Deblurring.
+We conduct deblurring experiments on RealDOF and DPDD datasets.
+To assess reconstruction fidelity and perceptual naturalness, we report LPIPS [83], DISTS [16], and no-reference metrics including CLIP-IQA [67], MANIQA [76], and MUSIQ [29].
+(ii) Bokeh Synthesis.
+We introduce LF-Bokeh, featuring $200$ images with diverse focus planes and aperture sizes synthesized from light-field captures [50, 14].
+In real-world scenarios, the absence of ground-truth depth maps and reference bokeh levels $K$ precludes the computation of accurate defocus maps, making the direct evaluation of fidelity and controllability challenging.
+To assess fidelity, following prior works [87, 43], we conduct a per-image binary search over $K$ and select the value that maximizes SSIM with the target, reporting LPIPS, DISTS, and CLIP-I [48].
+To quantify controllability under these constraints, we adopt the LVCorr metric following [80, 19, 69].
+Specifically, for images generated from the same all-in-focus input with a fixed focus plane across varying bokeh levels $K$, we compute the Pearson correlation coefficient (CorrCoef) between these $K$ values and the Laplacian variance trend, thereby reflecting the model’s accuracy in rendering the physical blur progression.
+(iii) Refocusing.
+The LF-Refocus dataset consists of $400$ source–target pairs from LF-Bokeh.
+Similar to bokeh synthesis, we employ an optimal $K$ search for each pair.
+We report LPIPS, DISTS, and CLIP-I for fidelity, alongside CLIP-IQA and MUSIQ to evaluate perceptual quality across varying depths of field.
+4.2 Comparison
+4.2.1 Defocus deblurring.
+As reported in Tab.
+2, DeblurNet outperforms competing methods across most referenced and non-referenced metrics.
+While our method yields slightly lower reference scores than Bokehlicious[57] on RealDOF[52], it consistently produces more visually appealing results and demonstrates greater robustness to diverse input conditions, such as grayscale photographs and low-quality, blurry images.
+Extensive visual comparison supporting this flexibility is provided in the supplementary material.
+Qualitatively, this perceptual superiority is evident in Fig.
+4.
+As shown in Fig.
+4(a), our approach achieves faithful recovery of text details with well-preserved edges.
+Furthermore, the scene in Fig.
+4(b) presents a particularly challenging case: while competing methods recover only a blurry dark stripe and introduce background distortions, DeblurNet reconstructs more consistent content and delivers compelling results.
+Table 4: Quantitative comparison on refocusing benchmark.
+Results on our LF-Refocus dataset.
+We compare our method against current state-of-the-art refocusing pipelines [69, 64].
+Optimal bokeh level $K$ is determined via per-image binary search following [87, 43].Learn2Refocus [64] takes a specific focal position as the input condition to synthesize the remaining images in the focal stack.
+To ensure coverage of the full range of focal planes, we perform inference twice separately: once conditioned on the minimum focal position (1) and once on the maximum (9).
+We then evaluate all outputs from both generated stacks and report the one yielding the highest SSIM.
+Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ CLIP-IQA $\uparrow$ MUSIQ $\uparrow$
+[…]
+[…]
+[…]
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Input GT DiffCamera Learn2Refocus GenRefocus
+[69] [64] (Ours)
+Figure 6: Qualitative comparison on the refocusing benchmark.
+Results on the LF-Refocus dataset with zoomed patches (blue and orange boxes) highlight the fidelity of focal and defocused regions.
+We observe that DiffCamera [69] struggles to reconstruct sharp structural details due to its strict input conditions and often assigns incorrect blurriness levels to background regions.
+Similarly, Learn2Refocus [64] fails to achieve accurate focus-plane placement, resulting in unintended blur of the target object.
+In contrast, our method achieves precise refocusing at the target plane while synthesizing photorealistic bokeh, closely matching the ground truth across diverse scenes.
+4.2.2 Bokeh synthesis.
+As presented in Tab.
+3, our approach outperforms all other methods across fidelity metrics.
+In terms of controllability, which measures how well the generated bokeh aligns with the conditioned bokeh level $K$, BokehMe [43] achieves near-perfect control owing to its hybrid design that incorporates a classical, physically motivated renderer.
+Nevertheless, our method outperforms all diffusion-based baselines, demonstrating that our learning strategy effectively preserves both realism and accuracy.
+Furthermore, it is worth noting the specific limitations of competing methods: DiffCamera [69] requires inputs with a fixed resolution and aspect ratio due to its reliance on explicit focus point conditioning, rendering it incompatible with tiling strategies; consequently, resizing operations severely degrade its controllability.
+Meanwhile, Bokeh Diffusion [19] relies on an inversion [61] process that not only prolongs inference time but also strips the image of its original high-frequency details, thereby lowering overall fidelity.
+The qualitative results shown in Fig.
+5.
+4.2.3 Refocusing.
+As shown in Tab.4, our approach achieves the best results on both fidelity and perceptual metrics, outperforming other baselines.
+Qualitatively, Fig.
+6 shows that competing methods tend to produce spatially ambiguous blur and fail to place the focal plane correctly, resulting in uniformly soft content.
+In contrast, our approach restores fine details at the intended focus while synthesizing physically consistent bokeh elsewhere.
+4.3 Ablation Study
+Table 5: Comparison of pipeline designs. Comparing direct refocusing with our proposed deblurring, then the bokeh synthesis approach on LF-Refocus.
 Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
-BokehMe [48] 0.1228 0.0744 0.9511
-Bokehlicious [60] 0.1799 0.1062 0.9304
-BokehDiff [89] 0.1708 0.0933 0.9192
-GenRefocus (Ours) 0.1047 0.0611 0.9570
-Benchmarks.
-We evaluate three tasks: (i) defocus deblurring on DPDD and RealDOF; (ii) bokeh synthesis on our new LF-Bokeh benchmark, comprising $200$ images with multiple focus planes synthesized from [53] and [15] light-field captures; and (iii) refocusing on LF-Refocus, built by pairing same-scene images within LF-Bokeh to form $400$ source–target pairs.
-Inference-time resolution and aspect ratio.
-We adopt a tiling strategy inspired by  [5] at evaluation time: we split the input image into overlapping patches and perform tile-wise denoising before blending them back together.
-As a result, all benchmarks are conducted at each dataset’s original aspect ratio and resolution, without additional resizing.
-4.2 Comparison
-Defocus deblurring.
-As stated on Tab.
-2, DeblurNet consistently outperforms all baselines across reference and no-reference metrics.
-On certain NR-IQA measures, it even surpasses the ground-truth.
-Qualitatively, Fig.
-4 (a) shows faithful recovery of the “NEW YORK” text with well-preserved edges.
-The scene in Fig.
-4 (b) is particularly challenging: competing methods recover only a blurry dark stripe and further distort the background, whereas DeblurNet reconstructs structure-consistent content and delivers visually compelling results without introducing artifacts or geometry drift.
-Table 4: Quantitative comparison on refocusing benchmark.
-Results on our LF-Refocus dataset (400 pairs with different focus planes).
-We compare our DeblurNet+BokehNet pipeline against combinations of existing all-in-focus estimation methods [56, 84] and bokeh synthesis methods [48, 89].
-Optimal bokeh level $K$ is determined via per-image binary search following [89].
-Our approach achieves superior performance across all metrics, demonstrating the benefit of joint optimization and semi-supervised training.
-AiF Estimation Bokeh Synthesis LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ NIQE $\downarrow$ CLIP-IQA $\uparrow$
-DRBNet [56] BokehMe [48] 0.1471 0.0937 0.9335 43.3087 6.5783 0.4460
-BokehDiff [89] 0.1848 0.1030 0.9310 48.0766 5.8195 0.4997
-Restormer [84] BokehMe [48] 0.1554 0.0963 0.9303 43.3188 7.1249 0.4782
-BokehDiff [89] 0.1860 0.1016 0.9320 48.0326 6.3216 0.5352
-DeblurNet (Ours) BokehNet (Ours) 0.1458 0.0850 0.9451 52.6954 5.3343 0.5763
-Bokeh synthesis.
-Following the BokehDiff protocol [89], we conduct a per-image binary search over the bokeh level $K$ and select the value that maximizes SSIM with the target.
-All other control signals are held fixed, ensuring a fair comparison in which each method operates at its best attainable setting.
-As summarized in Tab.
-3, our approach attains the highest scores on both fidelity and perceptual metrics, outperforming all baselines.
-The qualitative results in Fig.
-5 show that our method synthesizes bokeh that more closely matches that of real photographs.
-In the zoomed-in patches, our outputs more faithfully reproduce the ground-truth bokeh than competing baselines.
-Other methods tend to naively translate the predicted defocus map into blur strength, so any error in the depth scale directly leads to over- or under-blurring.
-In contrast, BokehNet, trained with our semi-supervised strategy, is more robust to such depth-scale inaccuracies and thus produces a more appropriate amount of blur that better matches the ground truth.
-Refocusing.
-We build four two-stage baselines by pairing strong AIF estimators (DRBNet [56], Restormer [84]) with bokeh synthesizers (BokehMe [48], BokehDiff [89]), yielding the $2\times 2$ combinations.
-For each method, we also follow the protocol of BokehDiff [89] to select the best bokeh level $K$ .
-As shown in Tab.4, our approach achieves the best results on both fidelity and perceptual metrics, outperforming all baseline pairings.
-Qualitatively, Fig.
-6 shows that competing methods tend to produce spatially ambiguous blur and fail to place the focal plane correctly, resulting in uniformly soft content and inaccurate foreground/background separation.
-In contrast, our approach restores fine details at the intended focus while synthesizing physically consistent bokeh elsewhere.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input GT DRBNet [56] DRBNet [56] Restormer [84] Restormer [84] GenRefocus
-+ BokehMe [48] + BokehDiff [89] + BokehMe [48] + BokehDiff [89] (Ours)
-Figure 6: Qualitative comparison on refocusing benchmark.
-Results on LF-Refocus comparing our end-to-end trained pipeline against baseline combinations of existing deblurring [56, 84] and bokeh synthesis [48, 89] methods.
-Zoomed patches (blue and orange boxes) highlight focal and defocused regions.
-Baselines struggle with accurate focus-plane placement and produce inconsistent or overly smooth blur.
-Our method achieves sharp refocusing at the target plane while synthesizing photorealistic bokeh, closely matching ground truth across diverse scenes.
-Table 5: Ablation on pipeline design.
-Comparing one-stage (direct refocusing) and two-stage (deblur then re-bokeh) on LF-Refocus.
-Both use the same backbone and synthetic data.
-Two-stage enables better depth control and independent semi-supervised learning per stage, yielding substantial improvements.
-Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
-One stage 0.1723 0.0976 0.9345
-Two stage (ours) 0.1458 0.0850 0.9461
-Table 6: Ablation on BokehNet training.
-Comparing synthetic-only vs.
-adding real unpaired bokeh images on LF-Bokeh.
-Second stage uses real bokeh with estimated parameters (Sec.
-3.2), initialized from synthetic weights.
-Real data enables learning authentic optical characteristics beyond simulators.
-Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
-Synthetic pair 0.1289 0.0738 0.9461
-Synthetic pair+ Real unpaired 0.1047 0.0611 0.9570
-4.3 Ablation Study
-One stage and two stage refocusing.
-We also evaluate a variant that takes an input image and a provided defocus map and directly produces a refocused output in a single pass.
-Concretely, we instantiate a one-stage model by reusing the BokehNet backbone and training it on the same synthetic pairs, then assess it on the refocusing benchmark.
-As reported in Tab.
-5, the one-stage variant lags the proposed two-stage pipeline by a substantial margin.
-We attribute this gap to two factors.
-First, control fidelity: the defocus map relies on depth estimated from a blurry input; without an all-in-focus view, depth quality degrades, leading to misaligned or biased blur fields and weakening refocusing accuracy.
-Second, data leverage: the one-stage formulation cannot separate the problem into defocus removal and bokeh synthesis, preventing us from injecting real-data supervision tailored to each subtask.
-Key idea and BokehNet training ablation.
-Our core idea is to use data generation to harvest real-world supervision for training BokehNet, leveraging real unpaired images.
-To quantify its impact, we evaluate two training settings without introducing additional architectural changes: (i) synthetic pair: BokehNet is trained purely on synthetic paired data; and (ii) synthetic pair + real unpaired: starting from (i), we further train with real unpaired images processed by our pipeline to produce usable supervision.
-On the bokeh-synthesis benchmark, the synthetic pair setting performs on par with a classical simulator baseline, while the synthetic pair + real unpaired setting delivers substantial improvements across both fidelity and perceptual metrics (Tab.
-6).
-Application.
-As illustrated in Fig.
-7, after we fine-tune BokehNet on our point-light 1K dataset, the model can synthesize bokeh with diverse, user-specified aperture shapes.
-Moreover, even though DeblurNet is trained without any text prompts, it can still leverage textual instructions at inference time to refine the all-in-focus output, as shown in Fig.
-8.
-This highlights a key advantage of generative defocus-deblurring models: they naturally support post-hoc, prompt-based correction and editing of the reconstructed content.
+Direct refocusing 0.1723 0.0976 0.9345
+Deblurring then bokeh synthesis (Ours) 0.1324 0.0759 0.9571
+Table 6: Ablation on BokehNet training datasets.
+Comparing combinations of datasets: (a) Synthetic data, (b) ITW dataset[19], and (c) LFDOF [52] and RealBokeh[57] on LF-Bokeh.
+The results show that combining all three yields the best optical characteristics.
+Datasets Metrics
+(a) Synthetic (b) ITW (c) LFDOF & RealBokeh LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
+- - 0.1289 0.0738 0.9461
+- 0.1156 0.0665 0.9529
+[…]
+[…]
+4.3.1 Direct refocusing vs. deblurring then bokeh synthesis.
+We evaluate a baseline that directly predicts a refocused image from a blurry input and a defocus map.
+Trained on identical synthetic data using the BokehNet backbone, this direct approach significantly underperforms our two-stage pipeline (Tab.
+5).
+We attribute this gap to two factors: (1) control fidelity: estimating depth from blurry inputs degrades defocus map accuracy, causing misaligned blur fields; and (2) data leverage: a direct formulation precludes injecting tailored real-data supervision into separate deblurring and synthesis subtasks.
+4.3.2 BokehNet training ablation.
+To quantify the impact of different training data distributions, we evaluate BokehNet training under various dataset combinations on LF-Bokeh.
+We consider three primary data sources: (a) synthetic data, (b) ITW dataset[19], and (c) LFDOF [52] and RealBokeh [57].
+As shown in Tab.
+6, a baseline trained purely on (a) provides a foundational performance but leaves room for improvement.
+Introducing either (b) or (c) to the synthetic baseline noticeably enhances performance.
+Ultimately, the comprehensive setting that combines all three sources yields the most substantial improvements, delivering the best optical characteristics and real-world generalization.
+4.3.3 Aperture shape control.
+As illustrated in Fig. 7, after we fine-tune BokehNet on our point-light 1K dataset, the model can synthesize bokeh with diverse, user-specified aperture shapes.
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Background point lights exhibit the specified aperture geometry while maintaining scene consistency.
-Figure 8: Text-guided deblurring.
-DeblurNet uses optional text prompts to guide reconstruction of severely blurred text.
-With prompts, we correct hallucinations (red: “DESION” $\rightarrow$ “DESIGN”) and recover accurate text (green: “yeurs” $\rightarrow$ “years”), matching GT.
-Without guidance, it hallucinates incorrect content.
+Figure 8: Failure case: transparent surfaces. With a transparent glass, focusing on the foreground may fail to correctly blur the background due to unreliable defocus maps.
+Figure 9: Failure case: hallucinated details. With severely blurred inputs, our model may hallucinate incorrect results.
 5 Conclusion
@@
 By decoupling into DeblurNet (deblurring) and BokehNet (bokeh synthesis), our approach handles arbitrary inputs with intuitive controls over focus plane, bokeh intensity, and aperture shape.
-Through semi-supervised training on synthetic and real unpaired bokeh data, GenRefocus learns authentic optical behavior beyond simulators.
-Experiments on DPDD, RealDOF, LF-Bokeh, and LF-Refocus show consistent improvements, enabling text-guided refinement and aperture-shape editing.
-Limitations and future work.
-GenRefocus relies on monocular depth estimation; severe failures result in mislocalized focal planes.
+Through training on various data sources, GenRefocus learns authentic optical behavior beyond simulators.
+Experiments on all benchmarks show consistent improvements, enabling aperture-shape editing.
+5.0.1 Limitations and future work.
+GenRefocus relies on the defocus map; Imperfections in the depth estimate can propagate to the defocus map and yield incorrect blur assignment.
+Transparent surfaces are particularly challenging, as monocular depth typically captures the depth of the front transparent surface while ignoring the background seen through it (see Fig.
+9).
+In cases of severe blur, our diffusion-based method may fail to recover fine details and instead hallucinate textures (see Fig.
+9), rendering it unsuitable for precision-critical fields such as scientific imaging or safety-sensitive applications.
 Complex aperture shapes require simulator-driven training data.
 Future work should generalize to richer aperture vocabularies, including user-drawn designs.
-Appendix A Appendix Overview
-In this appendix, we provide additional details on the training datasets used in our framework, the construction pipeline of our PointLight-1K dataset, a comparison with the VLM model, and visual comparisons from the BokehNet training ablation study.
-We also present failure cases and additional in-the-wild results on the accompanying HTML page.
-Figure 9: Comparison with VLM model. Qualitative refocusing results comparing our Generative Refocusing framework with Gemini 3 Nano Banana Pro [20] given prompt ”focus on the man on the right”.
-Appendix B Visual Comparison for BokehNet Training Ablation
-In the main paper, we compare different training strategies for BokehNet.
-As illustrated in Fig.
-10, under a foreground-bokeh setting with comparable bokeh strength in both rows, the model in (a), which is trained only on synthetic pairs, inherits artifacts from the simulator [48] and fails to correctly blur fine structures such as hair along object boundaries.
-In contrast, the model in (b), which is additionally trained with unpaired real bokeh data, produces smoother and more realistic blur around these edges.
-Appendix C Training Datasets
-We summarize the datasets used to train DeblurNet and BokehNet.
-C.1 DeblurNet
-For training DeblurNet, we use the DPDD [1] dataset and adopt all images from its official training split.
-This dataset provides paired defocused and all-in-focus images, which we use to supervise the deblurring stage.
-We further leverage the RealBokeh_3MP [60] dataset, which contains images captured with relatively small apertures and thus does not strictly provide all-in-focus views.
-To obtain images that are sufficiently sharp for our purposes, we compute the Laplacian variance of each image as a simple focus measure and rank the dataset accordingly.
-We then retain the top $3000$ images with the highest Laplacian variance as additional training data for DeblurNet.
-C.2 BokehNet
-For training BokehNet, we require diverse all-in-focus inputs.
-We therefore draw all-in-focus images from the [82] and [30] collections.
-As in the DeblurNet stage, we compute the Laplacian variance for each candidate image and filter out blurry examples.
-After this filtering, we obtain a pool of approximately 1.71.7k all-in-focus images for simulating bokeh images.
-The handling of the real bokeh datasets BokehDiffusion [19] and RealBokeh_3MP [60] for bokeh synthesis is already described in the main paper.
-Appendix D PointLight-1K Collection Pipeline
+Appendix 0.A Overview
+In the supplementary material, we provide additional details regarding the training datasets used in our framework and the construction pipeline of PointLight-1K dataset.
+We also present extended comparisons with DiffCamera [69], Bokehlicious [57], and Vision-Language Models.
+Furthermore, we provide qualitative results demonstrating the controllability of our approach, along with further analysis on ITW dataset [19] and a demonstration of the pre-deblur module’s effectiveness.
+We also present additional in-the-wild results on the accompanying HTML page.
+Appendix 0.B Additional Training Details
+In this section, we detail how we process datasets for training DeblurNet and BokehNet.
+0.B.1 DeblurNet Training Data
+DeblurNet is trained using a combination of paired defocus-blur data and high-quality sharp images.
+Specifically, we adopt all images from the official training split of the DPDD [1] dataset, which provides supervised pairs of defocused and all-in-focus images.
+To supplement this, we incorporate data from RealBokeh_3MP [57].
+Since this dataset contains images with varying degrees of focus, we apply a quality filter by computing the Laplacian variance of each image as a focus measure.
+By ranking the dataset based on this metric, we retain the top 3,000 sharpest images to serve as additional supervision for the deblurring task.
+0.B.2 BokehNet Training Data
+Training BokehNet requires diverse, high-resolution all-in-focus inputs to simulate realistic bokeh effects.
+We draw candidate images from [80] and the EBB [27] collections, and similar to the DeblurNet stage, we use Laplacian variance to filter out blurry examples.
+This yields a refined pool of approximately 1.7K sharp images.
+Furthermore, we utilize a comprehensive dataset of 26K real bokeh images.
+This collection comprises 13K previously filtered and verified images from the ITW dataset [19], alongside 13K images newly curated for this work.
+The newly collected data consists of focus-consistent series captured with varying apertures, containing 2 to 4 images per set.
+To establish accurate ground truth for these series, we first conducted a manual verification and refinement process on the in-focus masks.
+This annotation step required 4 to 8 seconds per image, amounting to approximately 8 hours of manual effort in total.
+Utilizing these masks, we then optimized the parameter $K$ using simulator [43].
+After obtaining the optimized $K$ values, we applied a Structural Similarity (SSIM) threshold to filter out sub-optimal results, ensuring that only reliable pairs are used for supervision.
+Appendix 0.C PointLight-1K Collection Pipeline
 We construct the PointLight-1K dataset to better model scenes with strong point-light bokeh. The pipeline consists of the following steps:
@@
 We refer to this dataset as PointLight-1K and use it as a specialized training set to pretrain and refine our learning strategy for controllable bokeh shape.
-Appendix E Comparison with VLM model
-We also compare our method with the recently released vision–language model Gemini-3 Nano Banana Pro [20], which explicitly supports text-driven refocusing.
+Table 7: Quantitative comparison on the DPDD dataset using $512\times 512$ central crops for the defocus deblurring task.
+Both methods are evaluated without any tiling strategy to ensure a strictly fair comparison of core architectural capacity.
+The best results are highlighted in bold.
+Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
+DiffCamera [69] 0.2833 0.1793 0.4602 0.3502 43.2997
+Ours 0.1285 0.1045 0.4637 0.3765 47.2070
+Appendix 0.D Additional Comparison with DiffCamera
+To ensure a strictly fair comparison of core model performance, we evaluate both methods under identical input conditions.
+Because DiffCamera [69] is constrained to its native $512\times 512$ training resolution and a fixed aspect ratio, we extract $512\times 512$ central crops from the DPDD benchmark.
+This guarantees that both models are evaluated at DiffCamera’s optimal setting.
+As shown in Table 7, our method consistently outperforms DiffCamera across all full-reference and no-reference metrics.
+Notably, our approach achieves a significant reduction in LPIPS (0.1285 vs.
+0.2833) and DISTS (0.1045 vs.
+0.1793).
+This comprehensive superiority under a strictly controlled baseline confirms our robust generative capacity and highlights the critical advantage of leveraging real-world data during training.
+Figure 10: Comparison with Bokehlicious [57].
+Results are shown for historical images (top three rows) and in-the-wild images (bottom three rows).
+Bokehlicious [57] either fails to remove the blur entirely or introduces severe structural artifacts.
+Conversely, our proposed method exhibits strong generalization capabilities, robustly handling diverse blur conditions to synthesize crisp and visually pleasing outputs.
+Appendix 0.E Qualitative comparison with Bokehlicious
+In the main paper, we state that our model demonstrates greater robustness to diverse input conditions.
+To further substantiate this claim, we provide additional qualitative comparisons with Bokehlicious [57] in Fig.
+10.
+A primary real-world application of defocus deblurring lies in restoring invaluable legacy photographs that can no longer be recaptured, as well as rescuing modern in-the-wild images suffering from accidental misfocus.
+As shown in the first three rows of Fig.
+10, we evaluate the models on historical images, including monochrome photographs.
+Under these conditions, Bokehlicious either yields negligible deblurring effects or introduces severe structural artifacts.
+Similarly, in the in-the-wild scenarios (the last three rows), Bokehlicious fails to effectively recover the sharp latent images and suffers from visual degradation.
+In contrast, our proposed method successfully restores details across all these highly diverse conditions.
+From aging black-and-white photos to everyday digital captures, our model consistently produces visually appealing results, thereby validating its strong generalization capabilities and practical applicability.
+Figure 11: Qualitative comparison with VLM model.
+Qualitative refocusing results comparing our Generative Refocusing framework with Gemini 3 Nano Banana Pro [20] given prompt "focus on the man on the right".
+Appendix 0.F Qualitative comparison with VLM model
+We also compare our method with the vision–language model Nano Banana Pro [20], which explicitly supports text-driven refocusing.
 For this model, we use the prompt “focus on the man on the right” to specify the desired focus point.
 As shown in Fig.
-9, Gemini-3 Nano Banana Pro can change the focus to some extent, but it also alters the facial expressions and appearances of people in both the foreground and the background.
+11, Nano Banana Pro can change the focus to some extent, but it also alters the facial expressions and appearances of people in both the foreground and the background.
 In contrast, our method preserves the original content while producing a realistic bokeh effect and accurate refocusing.
-Figure 10: Visual Comparison for BokehNet Training Ablation.
-Appendix F Failure Cases
-We observe two main types of unsatisfactory refocusing results: (1) incorrect hallucinated details and (2) inaccurate bokeh blur.
-For the first type, even though DeblurNet can generate visually plausible high-frequency details, in cases of severe input blur it may hallucinate incorrect detail (see Fig.
-11).
-For the second type, the bokeh blur can be imperfect because our defocus map is derived from monocular depth estimator [7].
-When the estimated depth scale is biased, the resulting defocus map may fail to fully cover the intended focus region or may assign too much or too little blur to certain areas.
-Although our semi-supervised training strategy for BokehNet mitigates these issues to some extent, residual discrepancies with respect to the ground truth can still be observed in challenging cases.
-Figure 11: Failure case. Due to the severely blurred input, the model hallucinates an incorrect time, changing 11:30 a.m. to 12:30 a.m.
+Figure 12: Qualitative evaluation of controllability.
+Refocusing results across different blur levels ($K\in\{0,5,10,15\}$).
+The blue boxes highlight specific text regions to illustrate the defocus effect.
+As $K$ increases, the blurriness of the scene increases monotonically, demonstrating our model’s precise control over the bokeh effect.
+Appendix 0.G Qualitative Results for Controllability
+In the main paper, we quantitatively demonstrated that our framework outperforms all other diffusion-based baselines in terms of controllability.
+In this section, we qualitatively evaluate this controllability by examining the influence of the blur parameter $K$ and the defocus map.
+As shown in figure 12, the blurriness of the scene increases monotonically as the given $K$ becomes larger.
+This monotonic transition effectively proves that our model achieves precise controllability.
+Figure 13: Camera model distribution. Percentage of images captured by the top 30 camera models in our dataset. Less frequent models are aggregated into the “Others” category.
+Figure 14: Focal length distribution. The percentage distribution of focal lengths across ITW dataset.
+Figure 15: Aperture (F-number) distribution. The percentage distribution of aperture settings across ITW dataset. Labels are presented as vertical fractions ($\frac{f}{N}$) for clarity.
+Figure 16: Pixel ratio distribution.
+The percentage distribution of the pixel-to-sensor ratio across ITW dataset.
+This ratio, defined as the image’s largest edge length divided by the physical sensor width ($\text{px}/\text{mm}$).
+Appendix 0.H ITW Dataset Analysis
+This dataset is utilized to estimate the image bokeh level $K$ via EXIF metadata.
+We provide further analysis regarding the dataset’s diversity.
+Crucially, all data consists of real optical bokeh images captured directly by digital cameras, rather than artificial defocus synthesized through computational aperture techniques.
+In addition to the hardware diversity represented by the various camera models (Fig.
+13), we analyze the distributions of the key parameters used in our $K$ estimation.
+Specifically, we provide the distributions for focal length (Fig.
+14), F-number (Fig.
+15), and pixel ratio (Fig.
+16).
+Figure 17: The effect of the pre-deblur module.
+Even when the pre-deblurred image is still distorted and blurry, using it as a condition yields higher-fidelity results (e.g., recovering “NEW YORK”) than removing this condition, which leads to distorted characters.
+Appendix 0.I The effect of the pre-deblur module.
+We introduce a variant that conditions DeblurNet on a pre-deblurred estimate.
+This provides a structural prior to disambiguate severe blur and improve fidelity.
+As shown in Fig.
+17, even if the pre-deblurred image remains distorted, it helps the model recover fine details (e.g., the text “NEW YORK”).